# A Hurewicz-type Theorem for the Dynamic Asymptotic Dimension with Applications to Coarse Geometry and Dynamics

### v4

@@ 1 Introduction
Fix $R>0$. Then there is $N$ such that $B_{e}^{R}(G)\subset G_{N}$. Notice that $d_{G}(G_{N}\cdot g,G_{N}\cdot h)$ is either 0 or >R>R. This follows since if $d_{G}(G_{N}\cdot g,G_{N}\cdot h)\leq R$ then the two orbits have a nontrivial intersection and therefore coincide. Since different $G_{N}$-orbits are >R>R apart, we can cover each one separately using $\text{asdim}G_{N}+1$ sets. For the second claim, find $N$ as before. Then if $x,y\in G/H_{j}$ are in separate $G_{N}$-orbits, their representatives in $G$ must be in separate $G_{N}$ orbits, and so $d_{G/{H_{j}}}(x,y)>R$. We can therefore similarly cover each orbit separately using $\text{asdim}\Box_{(H_{j}\cap G_{N})}G_{N}+1$ sets. ∎
@@
Suppose $\Gamma$ is a countable group with subgroup $N<\Gamma$ and recall we have equipped $\Gamma$ with a right-invariant, proper metric (see 2.2). Suppose further that $\Gamma/N$ is a quotient of $\Gamma$ with the metric induced from $\Gamma$. Consider the action $\Gamma\curvearrowright\Gamma/N$ on left cosets by left multiplication (this is an action even if $N$ is not normal). Then an $S$-bounded subset has diameter at most $\text{diam}(S)$ in $\Gamma$. Moreover, a subset $A\subset\Gamma/N$ with diameter at most $R$ is $B_{e}^{R}(\Gamma)$-bounded (recall that quotients do not increase the growth of a group). In particular, this implies a $(d,B_{e}^{r}(\Gamma),S)$-cover for $\Gamma\curvearrowright\Gamma/N$ is a $(d,r,\text{diam}(S))$-cover of $\Gamma/N$; and a $(d,r,R)$-cover of $\Gamma/N$ is a $(d,B_{e}^{r}(\Gamma),B_{e}^{R}(\Gamma))$-cover for $\Gamma\curvearrowright\Gamma/N$.
@@ 3 Odometers and Box Spaces
Let $\Gamma$ be a countable group and $(N_{i})$ ($i\in I$) a countable collection of finite-index subgroups of $\Gamma$ directed by inclusion. Then $\text{DAD}_{free}(\Gamma\curvearrowright\Gamma_{(N_{i})})\leq\text{asdim}\Box_{(N_{i})}\Gamma$.
@@
Suppose $\text{asdim}\Box_{(N_{i})}\Gamma\leq d$ and fix a finite subset $F\in\Gamma$. Find $R>0$ so that $d_{\Gamma}(\gamma\cdot x,x)<R$ for all $\gamma\in F$ and $x\in\Gamma$, hence $d_{\Gamma/N_{i}}(\gamma\cdot x,x)<R$ for all $\gamma\in F$ and all $i$. Suppose $\text{asdim}\Box_{(N_{i})}\Gamma\leq d$. Then $\text{asdim}(\Gamma/N_{i})\leq d$ uniformly. Let $\mathcal{U}^{i}=\{U^{i}_{0},\ldots,U^{i}_{n}\}$ be a $(d,R,M_{R})$-cover of $\Gamma/{N_{i}}$. Then by 2.16 each $\mathcal{U}^{i}$ is also a $(d,F,B_{e}^{M_{R}+1}(\Gamma/N_{i}))$-cover for $\Gamma\curvearrowright\Gamma/{N_{i}}$ (and since the cardinality of balls in finite quotients of $\Gamma$ is bounded by the cardinality of balls in $\Gamma$, this bound is independent of $i$). If $K=\ker(\Gamma\to\widehat{\Gamma}_{(N_{i})})$, then $\Gamma/K\curvearrowright\widehat{\Gamma}_{(N_{i})}$ freely and is orbit equivalent to the original $\Gamma$-odometer. Find $i_{F}$ so that the quotient $\Gamma/K\to\Gamma/N_{i_{F}}$ induces a map $C_{F}(\Gamma/K)\to C_{F}(\Gamma/N_{i_{F}})$ which is an isometry on balls of radius $\leq M_{R}+2$ (this is possible, for isntance, by [boxspacesDT, Proposition 2.1]). Then pulling back $\mathcal{U}^{i_{F}}$ by the projection $p_{i_{F}}:\widehat{\Gamma}_{(N_{i})}\to\Gamma/N_{i_{F}}$ gives a $(d,F,B_{e}^{M_{R}+1}(\Gamma/N_{i_{F}}))$-cover for $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$. ∎
@@
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite-index subgroups which is directed by inclusion (so $i\in\mathbb{N}$ and $N_{i+1}\subset N_{i}$). Let $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ be the odometer action induced by left multiplication on each finite quotient. If $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})\leq d$ then $\text{asdim}(\Box_{(N_{i})}\Gamma)\leq d$.
@@
Fix a proper, left-invariant metric on $\Gamma$. Let $R\geq 0$ be given and set $F=B_{e}^{R}(\Gamma)$. Let $\mathcal{V}=\{V_{0},\ldots,V_{d}\}$ be an open $(d,F,S)$-cover for $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$. Observe that for each $i$, the distance between the $F$-components of $p_{i}(V_{j})$ is >R>R according to the metric on $\Gamma/N_{i}$.
Now by 2.8 we can assume each $V_{j}$ is clopen and hence by 3.2 is the pullback of $T\subset\Gamma/N_{i}$ for some $i$. Find $i_{0}$ such that all of the $V_{j}$ are pullbacks of subsets of $\Gamma/N_{i_{0}}$. Then if $i\geq i_{0}$, $p_{i}(x)\in p_{i}(V_{j})$, and $f\cdot p_{i}(x)\in p_{i}(V_{j})$, then $x,f\cdot x\in V_{j}$. Thus, each $F$-component of $p_{i}(V_{j})$ is $S$-bounded for all $i\geq i_{0}$; so by 2.16 the diameter (as a subset of $\Box_{(N_{i})}\Gamma$) of each $R$-component of $p_{i}(V_{j})\subset\Gamma/N_{i}$ is bounded by $\text{diam}(S)$ for $i\geq i_{0}$. This gives a $(d,R,\text{diam}(S))$-cover of each $\Gamma/N_{i}$ for all $i\geq i_{0}$. Since $(N_{i})$ is a sequence, the set $\{i\mid i<i_{0}\}$ is finite, and so the diameter of $\Gamma/N_{i}$ for $i<i_{0}$ is uniformly bounded. We have therefore shown $\text{asdim}(\Gamma/N_{i})\leq d$ uniformly. The last part then follows from 2.14. ∎
@@
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite index subgroups which is directed by inclusion. Suppose $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ by left multiplication. Then $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\text{asdim}\Box_{(N_{i})}\Gamma$. ∎
@@
Let $\Gamma$ be a countable, residually finite group and $(N_{i})$ a sequence of finite index subgroups which is also a filtration of $\Gamma$ (see 2.11). Then $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\text{asdim}\Box_{(N_{i})}\Gamma$. ∎
@@
Let $\Gamma$ be a finitely generated, residually finite group and $(N_{i})$ a sequence of finite index subgroups which is also a filtration of $\Gamma$. Then either $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\infty$ or $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\text{asdim}\Gamma$. The dimension is finite if $\Gamma$ has polynomial growth (equivalently is virtually nilpotent). ∎
@@ 4 Product Theorem
Let $\Gamma\curvearrowright X$ be an action on a compact space by homeomorphisms and $A\subseteq X$. If $k\geq n+1\geq 1$, an $(n,k)$-dimensional control function for $A$ (with respect to $\Gamma\curvearrowright X$) is a function $\mathcal{D}_{A}:\mathcal{P}_{\text{fs}}(\Gamma)\to\mathcal{P}_{\text{fs}}(\Gamma)$ such that for any $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ there is a $(k-1,F,\mathcal{D}_{A}(F))$-cover $\mathcal{U}=\{U_{0},\ldots,U_{k-1}\}$ of $A$ (with respect to $\Gamma\curvearrowright X$) by sets open in the subspace topology on $A$ such that each $x\in A$ belongs to at least knk-n elements of $\mathcal{U}$ (equivalently $\cup_{i\in T}U_{i}=A$ for every $T\subset\{0,\ldots,k-1\}$ consisting of $n+1$ elements).
@@
(the trick) Let $\Gamma\curvearrowright X$ be an action on a compact metric space by homeomorphisms and $A\subset X$ closed.
@@
    Suppose $F,S\in\mathcal{P}_{\text{fs}}(\Gamma)$, $n\geq 0$, $k\geq n+1$, and $\mathcal{U}=\{U_{0},\ldots,U_{k-1}\}$ is a (k1,F,S)(k-1,F,S)-cover for $A$ by sets open in the subspace topology on $A$ such that each $x$ belongs to $\geq k-n$ elements. Then there is a (k1,F,S)(k-1,F,S)-cover by sets which are the closures of open sets (so closed in $X$) and still have the property that each $x$ belongs to $\geq k-n$ elements.
@@
    Suppose the $F$-components of $A$ are $S$-bounded. Then $A$ is contained in a set which is open in $X$ and whose $F$-components are $S$-bounded.
@@
We can assume $S\in\mathcal{P}_{fs}(\Gamma)$. To prove $(2)$, suppose not, that is, suppose for every $n$, the $F$ components of $N_{1/n}(U)$ are not $S$-bounded. So for every $n$, we have an $F$-chain $x^{(n)}_{0},\ldots,x^{(n)}_{K_{n}}$ in $N_{1/n}(U)$ with $x^{(n)}_{0},\ldots,x^{(n)}_{K_{n}-1}\in S\cdot x_{0}$ but $x_{K_{n}}=\gamma_{n}\cdot x^{(n)}_{0}$ for $\gamma_{n}=f_{n}s_{n}$ with $s_{n}\in S$, $f_{n}\in F$, and $\gamma_{n}\notin S$. Since there are finitely-many $\gamma\in\Gamma$ with the property that $\gamma=fs$ for $s\in S$ and $f\in F$, we must have $\gamma_{n}=\lambda$ for infinitely-many $n$. Furthermore, there are finitely-many sequences $f_{1},\ldots,f_{K_{n}}$ ($f_{k}\in F$) such that $f_{k}\cdots f_{1}\in S$ for all $1\leq k<K_{n}$ but $f_{K_{n}}\cdots f_{1}=\lambda\notin S$, and so for infinitely-many $n$ we have an $F$-chain $x^{(n)}_{0},\ldots,x^{(n)}_{K}$ in $N_{1/n}(U)$ and $f_{1},\ldots,f_{K}\in F$ such that $f_{k+1}\cdot x^{(n)}_{k}=x^{(n)}_{k+1}$ for $0\leq k<K$, and also $x_{K}=\lambda\cdot x_{0}$ with $\lambda\notin S$. Since $X$ is compact, we have a subsequence $x_{0}^{(n_{j})}$ converging to a point $y_{0}\in X$, and so $f_{k}\ldots f_{1}\cdot x_{0}^{(n_{j})}=x_{k}^{(n_{j})}$ converges to $f_{k}\ldots f_{1}\cdot y_{0}$ for all $1\leq k\leq K$ by continuity. Since $x_{k}^{(n_{j})}\in N_{1/n_{j}}(U)$ and $U$ is closed in $X$, $x_{k}^{(n_{j})}$ converges to a point in $U$. Hence, we have an $F$-chain in $U$ which is not $S$-bounded. ∎
@@
(compare to [Brodskiy2006AHT, Theorem 2.4]) Suppose $A\subset X$ is closed. Suppose $\mathcal{D}_{A}^{(n+1)}$ is an $n$-dimensional control function for $A$. Define $\{\mathcal{D}_{A}^{(i)}\}_{i\geq n+1}$ inductively on powers of $F$ by $\mathcal{D}_{A}^{(i+1)}(F^{r}):=F^{r}\cdot\mathcal{D}^{(i)}_{A}(F^{3r})\cdot F^{r}$; then for each $k$, there is an open $(k-1,F^{r},\mathcal{D}^{(k)}_{A}(F^{r}))$-cover of $A$ such that each $x\in A$ is covered by at least knk-n elements of the cover. (We say it this way since the functions $\mathcal{D}_{A}^{(k)}$ are not defined on all elements of $\mathcal{P}_{\text{fs}}(\Gamma)$ unless $\Gamma$ is finitely generated, so we can’t quite call them control functions).
@@
The case $k=n+1$ is obvious. Suppose the result holds for some larger $k$. Let $\mathcal{U}=\{U_{1},\ldots,U_{k}\}$ be an open $(k-1,F^{3r},\mathcal{D}_{A}^{(k)}(F^{3r}))$-cover for $A$ with all $x\in A$ in at least knk-n elements of $\mathcal{U}$. Using (1) from the trick (see 4.3), we can replace the sets $U_{i}$ by the closures of open sets, $V_{i}$, which form a cover of $A$ with the same properties. Define $U_{i}^{\prime}=(F^{r}\cdot U_{i})\cap A$ (these are closed sets). Then $\mathcal{U}^{\prime}=\{U_{1}^{\prime},\ldots,U_{k}^{\prime}\}$ is a closed $(k-1,F^{r},F^{r}\cdot\mathcal{D}_{A}^{(k)}(F^{3r})\cdot F^{r})$-cover for $A$. Notice that $F^{r}\cdot\mathcal{D}_{A}^{(k)}(F^{3r})\cdot F^{r}=\mathcal{D}_{A}^{(k+1)}(F^{r})$.
@@
Fix $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ and $H\in\mathcal{P}_{\text{fs}}(\Lambda)$. Suppose $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ have dimension at most $m$ and at most $n$, respectively, and let $\mathcal{D}_{X}$ and $\mathcal{D}_{Y}$ be respective control functions. Let $k=m+n+1$. Use 4.4 to produce the functions $\mathcal{D}^{(k)}_{X}$ and $\mathcal{D}^{(k)}_{Y}$. So there is an open $(k-1,F,\mathcal{D}^{(k)}_{\Gamma\curvearrowright X}(F))$-cover $\mathcal{U}=\{U_{i}\}_{i=0}^{k-1}$ of $X$ and an open $(k-1,H,\mathcal{D}^{(k)}_{\Lambda\curvearrowright Y}(H))$-cover $\mathcal{V}=\{V_{i}\}_{i=0}^{k-1}$ of $Y$ such that each $x\in X$ is in at least kmk-m elements of $\mathcal{U}$ and each $y\in Y$ is in at least knk-n elements of $\mathcal{V}$. Since km+kn=k+1k-m+k-n=k+1, the family $\{U_{i}\times V_{i}\}_{i=0}^{k-1}$ is a cover of $X\times Y$. Moreover, it is a $(F\times H,\mathcal{D}^{(k)}_{\Gamma\curvearrowright X}(F)\cdot\mathcal{D}^{(k)}_{\Lambda\curvearrowright Y}(H))$-cover for $\Gamma\times\Lambda\curvearrowright X\times Y$. ∎

### v6

@@ 1 Introduction
Fix $R>0$. Then there is $N$ such that $B_{e}^{R}(G)\subset G_{N}$. Notice that $d_{G}(G_{N}\cdot g,G_{N}\cdot h)$ is either 0 or >R>R. This follows since if $d_{G}(G_{N}\cdot g,G_{N}\cdot h)\leq R$ then the two orbits have a nontrivial intersection and therefore coincide. Since different $G_{N}$-orbits are >R>R apart, we can cover each one separately using $\operatorname{\text{asdim}}G_{N}+1$ sets. For the second claim, find $N$ as before. Then if $x,y\in G/H_{j}$ are in separate $G_{N}$-orbits, their representatives in $G$ must be in separate $G_{N}$ orbits, and so $d_{G/{H_{j}}}(x,y)>R$. We can therefore similarly cover each orbit separately using $\operatorname{\text{asdim}}\Box_{(H_{j}\cap G_{N})}G_{N}+1$ sets. ∎
@@
Suppose $\Gamma$ is a countable group with subgroup $N<\Gamma$ and recall we have equipped $\Gamma$ with a right-invariant, proper metric (see 2.2). Suppose further that $\Gamma/N$ is a quotient of $\Gamma$ with the metric induced from $\Gamma$. Consider the action $\Gamma\curvearrowright\Gamma/N$ on left cosets by left multiplication (this is an action even if $N$ is not normal). Then an $S$-bounded subset has diameter at most $\text{diam}(S)$ in $\Gamma$. Moreover, a subset $A\subset\Gamma/N$ with diameter at most $R$ is $B_{e}^{R}(\Gamma)$-bounded (recall that quotients do not increase the growth of a group). In particular, this implies a $(d,B_{e}^{r}(\Gamma),S)$-cover of $\Gamma/N$ is a $(d,r,\text{diam}(S))$-cover of $\Gamma/N$; and a $(d,r,R)$-cover of $\Gamma/N$ is a $(d,B_{e}^{r}(\Gamma),B_{e}^{R}(\Gamma))$-cover of $\Gamma/N$.
@@ 3 Odometers and Box Spaces
Let $\Gamma$ be a countable group and $(N_{i})$ ($i\in I$) a countable collection of finite-index subgroups of $\Gamma$ directed by inclusion. Then $\text{DAD}_{free}(\Gamma\curvearrowright\Gamma_{(N_{i})})\leq\operatorname{\text{asdim}}\Box_{(N_{i})}\Gamma$.
@@
Suppose $\operatorname{\text{asdim}}\Box_{(N_{i})}\Gamma\leq d$ and fix a finite subset $F\in\Gamma$. Find $R>0$ so that $d_{\Gamma}(\gamma\cdot x,x)<R$ for all $\gamma\in F$ and $x\in\Gamma$, hence $d_{\Gamma/N_{i}}(\gamma\cdot x,x)<R$ for all $\gamma\in F$ and all $i$. Suppose $\operatorname{\text{asdim}}\Box_{(N_{i})}\Gamma\leq d$. Then $\operatorname{\text{asdim}}(\Gamma/N_{i})\leq d$ uniformly. Let $\mathcal{U}^{i}=\{U^{i}_{0},\ldots,U^{i}_{n}\}$ be a $(d,R,M_{R})$-cover of $\Gamma/{N_{i}}$. Then by 2.16 each $\mathcal{U}^{i}$ is also a $(d,F,B_{e}^{M_{R}+1}(\Gamma/N_{i}))$-cover of $\Gamma/{N_{i}}$ (and since the cardinality of balls in finite quotients of $\Gamma$ is bounded by the cardinality of balls in $\Gamma$, this bound is independent of $i$). If $K=\ker(\Gamma\to\widehat{\Gamma}_{(N_{i})})$, then $\Gamma/K\curvearrowright\widehat{\Gamma}_{(N_{i})}$ freely and is orbit equivalent to the original $\Gamma$-odometer. Find $i_{F}$ so that the quotient $\Gamma/K\to\Gamma/N_{i_{F}}$ induces a map $C_{F}(\Gamma/K)\to C_{F}(\Gamma/N_{i_{F}})$ which is an isometry on balls of radius $\leq M_{R}+2$ (this is possible, for isntance, by [boxspacesDT, Proposition 2.1]). Then pulling back $\mathcal{U}^{i_{F}}$ by the projection $p_{i_{F}}:\widehat{\Gamma}_{(N_{i})}\to\Gamma/N_{i_{F}}$ gives a $(d,F,B_{e}^{M_{R}+1}(\Gamma/N_{i_{F}}))$-cover of $\widehat{\Gamma}_{(N_{i})}$. ∎
@@
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite-index subgroups which is directed by inclusion (so $i\in\mathbb{N}$ and $N_{i+1}\subset N_{i}$). Let $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ be the odometer action induced by left multiplication on each finite quotient. If $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})\leq d$ then $\operatorname{\text{asdim}}(\Box_{(N_{i})}\Gamma)\leq d$.
@@
Fix a proper, left-invariant metric on $\Gamma$. Let $R\geq 0$ be given and set $F=B_{e}^{R}(\Gamma)$. Let $\mathcal{V}=\{V_{0},\ldots,V_{d}\}$ be an open $(d,F,S)$-cover of $\widehat{\Gamma}_{(N_{i})}$. Observe that for each $i$, the distance between the $F$-components of $p_{i}(V_{j})$ is >R>R according to the metric on $\Gamma/N_{i}$.
Now by 2.8 we can assume each $V_{j}$ is clopen and hence by 3.2 is the pullback of $T\subset\Gamma/N_{i}$ for some $i$. Find $i_{0}$ such that all of the $V_{j}$ are pullbacks of subsets of $\Gamma/N_{i_{0}}$. Then if $i\geq i_{0}$, $p_{i}(x)\in p_{i}(V_{j})$, and $f\cdot p_{i}(x)\in p_{i}(V_{j})$, then $x,f\cdot x\in V_{j}$. Thus, each $F$-component of $p_{i}(V_{j})$ is $S$-bounded for all $i\geq i_{0}$; so by 2.16 the diameter (as a subset of $\Box_{(N_{i})}\Gamma$) of each $R$-component of $p_{i}(V_{j})\subset\Gamma/N_{i}$ is bounded by $\text{diam}(S)$ for $i\geq i_{0}$. This gives a $(d,R,\text{diam}(S))$-cover of each $\Gamma/N_{i}$ for all $i\geq i_{0}$. Since $(N_{i})$ is a sequence, the set $\{i\mid i<i_{0}\}$ is finite, and so the diameter of $\Gamma/N_{i}$ for $i<i_{0}$ is uniformly bounded. We have therefore shown $\operatorname{\text{asdim}}(\Gamma/N_{i})\leq d$ uniformly. The last part then follows from 2.14. ∎
@@
Let $\Gamma$ be a countable group and $(N_{i})$ a sequence of finite index subgroups which is directed by inclusion. Suppose $\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})}$ by left multiplication. Then $\text{DAD}_{free}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\operatorname{\text{asdim}}\Box_{(N_{i})}\Gamma$. ∎
@@
Let $\Gamma$ be a countable, residually finite group and $(N_{i})$ a sequence of finite index subgroups which is also a filtration of $\Gamma$ (see 2.11). Then $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\operatorname{\text{asdim}}\Box_{(N_{i})}\Gamma$. ∎
@@
Let $\Gamma$ be a finitely generated, residually finite group and $(N_{i})$ a sequence of finite index subgroups which is also a filtration of $\Gamma$. Then either $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\infty$ or $\text{DAD}(\Gamma\curvearrowright\widehat{\Gamma}_{(N_{i})})=\operatorname{\text{asdim}}\Gamma$. The dimension is finite if $\Gamma$ has polynomial growth (equivalently is virtually nilpotent). ∎
@@ 4 Product Theorem
Let $\Gamma\curvearrowright X$ be an action on a compact metric space by homeomorphisms and $A\subseteq X$. If $k\geq n+1\geq 1$, an $(n,k)$-dimensional control function for $A$ (with respect to $\Gamma\curvearrowright X$) is a function $\mathcal{D}_{A}:\mathcal{P}_{\text{fs}}(\Gamma)\to\mathcal{P}_{\text{fs}}(\Gamma)$ such that for any $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ there is a $(k-1,F,\mathcal{D}_{A}(F))$-cover $\mathcal{U}=\{U_{0},\ldots,U_{k-1}\}$ of $A$ (with respect to $\Gamma\curvearrowright X$) by sets open in the subspace topology on $A$ such that each $x\in A$ belongs to at least knk-n elements of $\mathcal{U}$ (equivalently $\cup_{i\in T}U_{i}=A$ for every $T\subset\{0,\ldots,k-1\}$ consisting of $n+1$ elements).
@@
(the trick) Let $\Gamma\curvearrowright X$ be an action on a compact metric space by homeomorphisms and $A\subset X$ closed. Let $F,S\in\mathcal{P}_{\text{fs}}(\Gamma)$.
@@
    If $n\geq 0$, $k\geq n+1$, and $\mathcal{U}=\{U_{0},\ldots,U_{k-1}\}$ is a (k1,F,S)(k-1,F,S)-cover of $A$ by sets open in the subspace topology on $A$ such that each $x$ belongs to $\geq k-n$ elements. Then there is a (k1,F,S)(k-1,F,S)-cover by sets which are the closures of open sets (so closed in $X$) and still have the property that each $x$ belongs to $\geq k-n$ elements.
@@
    If the $F$-components of $A$ are $S$-bounded, then $A$ is contained in a set which is open in $X$ and whose $F$-components are $S$-bounded.
@@
To prove $(2)$, suppose not, that is, suppose for every $n$ that the $F$ components of $N_{1/n}(A)$ are not $S$-bounded. So for every $n$, we have an $F$-chain $x^{(n)}_{0},\ldots,x^{(n)}_{K_{n}}$ in $N_{1/n}(A)$ with $x^{(n)}_{0},\ldots,x^{(n)}_{K_{n}-1}\in S\cdot x_{0}$ but $x_{K_{n}}=\gamma_{n}\cdot x^{(n)}_{0}$ for $\gamma_{n}=f_{n}s_{n}$ with $s_{n}\in S$, $f_{n}\in F$, and $\gamma_{n}\notin S$. Since there are finitely-many $\gamma\in\Gamma$ with the property that $\gamma=fs$ for $s\in S$ and $f\in F$, we must have $\gamma_{n}=\lambda$ for infinitely-many $n$. Furthermore, there are finitely-many sequences $f_{1},\ldots,f_{K_{n}}$ ($f_{k}\in F$) such that $f_{k}\cdots f_{1}\in S$ for all $1\leq k<K_{n}$ but $f_{K_{n}}\cdots f_{1}=\lambda\notin S$, and so for infinitely-many $n$ we have an $F$-chain $x^{(n)}_{0},\ldots,x^{(n)}_{K}$ in $N_{1/n}(A)$ and $f_{1},\ldots,f_{K}\in F$ such that $f_{k+1}\cdot x^{(n)}_{k}=x^{(n)}_{k+1}$ for $0\leq k<K$, and also $x_{K}=\lambda\cdot x_{0}$ with $\lambda\notin S$. Since $X$ is compact, we have a subsequence $x_{0}^{(n_{j})}$ converging to a point $y_{0}\in X$, and so $f_{k}\ldots f_{1}\cdot x_{0}^{(n_{j})}=x_{k}^{(n_{j})}$ converges to $f_{k}\ldots f_{1}\cdot y_{0}$ for all $1\leq k\leq K$ by continuity. Since $x_{k}^{(n_{j})}\in N_{1/n_{j}}(A)$ and $A$ is closed in $X$, $x_{k}^{(n_{j})}$ converges to a point in $A$. Hence, we have an $F$-chain in $A$ which is not $S$-bounded. ∎
@@
(compare to [Brodskiy2006AHT, Theorem 2.4]) Suppose $A\subset X$ is closed and that $\mathcal{D}_{A}^{(n+1)}$ is an $n$-dimensional control function for $A$. Define $\{\mathcal{D}_{A}^{(i)}\}_{i\geq n+1}$ inductively on powers of $F$ by $\mathcal{D}_{A}^{(i+1)}(F^{r}):=F^{r}\cdot\mathcal{D}^{(i)}_{A}(F^{3r})\cdot F^{r}$; then for each $k$, there is an open $(k-1,F^{r},\mathcal{D}^{(k)}_{A}(F^{r}))$-cover of $A$ such that each $x\in A$ is covered by at least knk-n elements of the cover. (We say it this way since the functions $\mathcal{D}_{A}^{(k)}$ are not defined on all elements of $\mathcal{P}_{\text{fs}}(\Gamma)$ unless $\Gamma$ is finitely generated, so we can’t quite call them control functions).
@@
The case $k=n+1$ is obvious. Suppose the result holds for some larger $k$. Let $\mathcal{U}=\{U_{1},\ldots,U_{k}\}$ be an open $(k-1,F^{3r},\mathcal{D}_{A}^{(k)}(F^{3r}))$-cover of $A$ with all $x\in A$ in at least knk-n elements of $\mathcal{U}$. Using (1) from the trick (see 4.3), we can replace the sets $U_{i}$ by the closures of open sets, $V_{i}$, which form a cover of $A$ with the same properties. Define $U_{i}^{\prime}=(F^{r}\cdot U_{i})\cap A$ (these are closed sets). Then $\mathcal{U}^{\prime}=\{U_{1}^{\prime},\ldots,U_{k}^{\prime}\}$ is a closed $(k-1,F^{r},F^{r}\cdot\mathcal{D}_{A}^{(k)}(F^{3r})\cdot F^{r})$-cover of $A$. Recall that $F^{r}\cdot\mathcal{D}_{A}^{(k)}(F^{3r})\cdot F^{r}=\mathcal{D}_{A}^{(k+1)}(F^{r})$ by definition.
@@
Fix $F\in\mathcal{P}_{\text{fs}}(\Gamma)$ and $H\in\mathcal{P}_{\text{fs}}(\Lambda)$. Suppose $\Gamma\curvearrowright X$ and $\Lambda\curvearrowright Y$ have dimension at most $m$ and at most $n$, respectively, and let $\mathcal{D}_{X}$ and $\mathcal{D}_{Y}$ be respective control functions. Let $k=m+n+1$. Use 4.4 to produce the functions $\mathcal{D}^{(k)}_{X}$ and $\mathcal{D}^{(k)}_{Y}$. So there is an open $(k-1,F,\mathcal{D}^{(k)}_{\Gamma\curvearrowright X}(F))$-cover $\mathcal{U}=\{U_{i}\}_{i=0}^{k-1}$ of $X$ and an open $(k-1,H,\mathcal{D}^{(k)}_{\Lambda\curvearrowright Y}(H))$-cover $\mathcal{V}=\{V_{i}\}_{i=0}^{k-1}$ of $Y$ such that each $x\in X$ is in at least kmk-m elements of $\mathcal{U}$ and each $y\in Y$ is in at least knk-n elements of $\mathcal{V}$. Since km+kn=k+1k-m+k-n=k+1, the family $\{U_{i}\times V_{i}\}_{i=0}^{k-1}$ is a cover of $X\times Y$. Moreover, it is a $(F\times H,\mathcal{D}^{(k)}_{\Gamma\curvearrowright X}(F)\cdot\mathcal{D}^{(k)}_{\Lambda\curvearrowright Y}(H))$-cover of $\Gamma\times\Lambda\curvearrowright X\times Y$. ∎